# AdaptNC: Adaptive Nonconformity Scores for Uncertainty-Aware Autonomous Systems in Dynamic Environments

### v1

@@ Abstract
Rigorous uncertainty quantification is essential for the safe deployment of autonomous systems in unconstrained environments. Conformal Prediction (CP) provides a distribution-free framework for this task, yet its standard formulations rely on exchangeability assumptions that are violated by the distribution shifts inherent in real-world robotics. Existing online CP methods maintain target coverage by adaptively scaling the conformal threshold, but typically employ a static nonconformity score function. We show that this fixed geometry leads to highly conservative, volume-inefficient prediction regions when environments undergo structural shifts. To address this, we propose AdaptNC, a framework for the joint online adaptation of both the nonconformity score parameters and the conformal threshold. AdaptNC leverages an adaptive reweighting scheme to optimize score functions, and introduces a replay buffer mechanism to mitigate the coverage instability that occurs during score transitions. We evaluate AdaptNC on diverse robotic benchmarks involving multi-agent policy changes, environmental changes and sensor degradation. Our results demonstrate that AdaptNC significantly reduces prediction region volume compared to state-of-the-art threshold-only baselines while maintaining target coverage levels.
Machine Learning, ICML
@@ 1 Introduction
Autonomous systems operating in real-world environments face a constant challenge: distribution shift. These shifts, driven by changing environmental conditions, system degradation, or complex multi-agent interactions, can severely degrade the performance of predictive models. For safety-critical applications such as autonomous driving and robotic navigation, identifying these shifts and providing robust, uncertainty-aware predictions is not just a performance requirement; it is a safety necessity (Lindemann et al., 2025, 2023; Yang et al., 2023; Tayal et al., 2025). Identifying this distribution shift and quantifying the uncertainty of predictors in these settings is a significant challenge. Conformal Prediction (CP) has emerged as a gold-standard framework for providing such reliability, offering distribution-free uncertainty quantification with formal coverage guarantees (Vovk et al., 2022). Although split conformal methods are widely used due to their simplicity and guarantees, they typically assume that the data distribution remains exchangeable with a held-out calibration set (Lei et al., 2018), an assumption which is violated under distribution shift.
Recent advances in Online Conformal Prediction have addressed this violation by adjusting the conformal quantile thresholds to maintain target coverage levels. This is commonly done through reweighting techniques (Tibshirani et al., 2019; Barber et al., 2023) or adaptive quantile adjustments (Gibbs & Candes, 2021; Gibbs & Candès, 2024; Zaffran et al., 2022). Although effective in achieving target coverage, these methods suffer from a critical limitation: they treat the nonconformity score function as fixed. The score function determines the geometry and efficiency of prediction regions. A static score function in a shifting environment often results in unnecessarily large conservative uncertainty bounds that lead to suboptimal control and decision-making. The choice and parameterization of these nonconformity scores significantly impact not only the efficiency (i.e., volume) of the prediction regions but also their informativeness for downstream decision-making tasks (Tumu et al., 2024).
@@
In this paper, we propose AdaptNC, a novel framework that enables the joint online adaptation of both the conformal threshold and the nonconformity score parameters. Unlike existing methods that only tune the threshold, AdaptNC optimizes the underlying score function to ensure that prediction regions remain tight and informative even as the data distribution evolves. To handle non-stationarity, we introduce an adaptive reweighting scheme that leverages DtACI expert weights to prioritize the most relevant historical observations during score optimization. To mitigate “coverage shock” (the sudden instability in coverage that occurs when a score function is modified) we employ a counterfactual replay mechanism to recalibrate the quantile threshold following a score update.
@@
Contributions: This paper addresses the challenge of constructing efficient, uncertainty-aware prediction regions under arbitrary distribution shift by contributing:
@@
    Online Score Adaptation: We present a method for the online optimization of nonconformity scores and their corresponding quantiles that minimizes prediction region volume while enabling adaptation to shifting distributions.
@@
    Dynamic Weighting and Replay Buffer: We develop a weighting mechanism for data based on distribution shift, and a replay mechanism to maintain coverage stability during score parameter transitions.
@@
    Empirical Validation on Robotic Benchmarks: We evaluate AdaptNC on robotic tasks involving structural shifts, including multi-agent policy changes and sensor degradation. We demonstrate that AdaptNC significantly reduces prediction volume compared to state-of-the-art online CP baselines while rigorously maintaining target coverage levels.
@@ 2 Problem Setup and Preliminaries
2.1 Online Regression Setting
We consider an online regression setting where, at each time step $t=1,2\ldots,T$, the learner observes a feature vector $X_{t}\in\mathcal{X}$, produces a point prediction $\hat{Y}_{t}=\{\hat{y}_{t+1},\ldots,\hat{y}_{t+\tau}\}\in\mathcal{Y}^{\tau}$ for outcomes over $\tau$ steps, and subsequently observes the true outcome $Y_{t}\in\mathcal{Y}^{\tau}$. A predictor $h:\mathcal{X}\rightarrow\mathcal{Y}^{\tau}$ is used to generate the point predictions, i.e., $\hat{Y}_{t}=h(X_{t})$. The joint distribution of $(X_{t},Y_{t})$ may change over time, reflecting distribution shift.
We seek to build a set where the true outcome $Y_{t}$ is above this set with probability $\alpha$. This is known as the miscoverage rate, with $\delta=1-\alpha$ being the coverage rate. The sets are known as prediction regions $\mathcal{C}_{t}(X_{t};\theta_{t},q_{\delta,\mathcal{D}_{t}})\subseteq\mathcal{Y}^{\tau}$ for each time $t$ with a target miscoverage rate of $\alpha$ (equivalently, coverage rate $1-\alpha$). The variable $q_{\delta,\mathcal{D}_{t}}$ is the $\delta$ quantile of the dataset $\mathcal{D}_{t}$. The prediction regions are constructed using a non-conformity score function $s(X_{t},y;\theta_{t})$ where $\theta_{t}$ are learnable parameters updated online. Specifically, the prediction region is defined as: $\mathcal{C}_{t}(X_{t};\theta_{t},q_{\delta,\mathcal{D}_{t}})\coloneqq\{y\in\mathcal{Y}^{\tau}:s(X_{t},y;\theta_{t})\leq q_{\delta,\mathcal{D}_{t}}\}$.
2.2 Desiderata
We pursue two main objectives in this online regression setting, coverage and efficiency, under distribution shift.
Coverage.
Due to distribution shift, we do not expect to meet this coverage target at all times. Instead, we seek the asymptotic property that our long-run miscoverage rate approaches $\alpha$, under the assumption that the distribution stabilizes after a period of shift. We define the error indicator $\operatorname{\textbf{err}}_{t}=\mathds{1}\{Y_{t}\notin\mathcal{C}_{t}(X_{t};\theta_{t},q_{\delta,\mathcal{D}_{t}})\}$, and state our desired coverage property as follows:
@@ 2.2 Desiderata
Efficiency
We would also like to ensure that the prediction regions $\mathcal{C}_{t}$ are small and informative to downstream algorithms. Thus, while maintaining the coverage guarantee above, our objective is to minimize the volume of the prediction set $\mathcal{C}_{t}$. We formulate the optimization problem:
$\min_{\theta_{t}}\quad$ $\text{Vol}(\mathcal{C}_{t}(X_{t+1};\theta_{t},q_{\delta,\mathcal{D}_{t}}))$ (2a)
s.t. $\Pr[Y_{t+1}\in\mathcal{C}_{t}(X_{t+1};\theta_{t},q_{\delta,\mathcal{D}_{t}})]\geq 1-\alpha.$ (2b)
@@ 2.3 Preliminaries
Conformal Prediction.
Conformal prediction (CP) provides distribution-free, finite-sample prediction regions which guarantee marginal coverage with high probability under the assumption of exchangeable data (Vovk et al., 2022; Lei et al., 2018). Split conformal prediction and its variants, like jackknife+ (Lei et al., 2018; Barber et al., 2020), extend this framework to modern machine learning train-test paradigms while reducing the computational cost of full conformal methods. Much of this literature treats the non-conformity score as fixed, with scores resembling residuals or norms for regression tasks, and focuses on how to calibrate the conformal quantile to ensure coverage. Our work targets the non-conformity score itself, in order to provide efficient and adaptive prediction regions. The guarantee that Conformal Prediction seeks to provide is marginal coverage, i.e., $P(Y_{t}\in\mathcal{C}_{t}(X_{t};\theta_{t},q_{\delta,\mathcal{D}_{t}}))\geq 1-\alpha$. This guarantee holds in finite samples under exchangeability, but is not guaranteed to work under distribution shift.
Maintaining Coverage under Distribution Shift.
In settings where distribution shift is present, the exchangeability assumption is violated, and conformal prediction methods may fail to maintain coverage. Our method builds on recent work that extends conformal prediction to settings with distribution shift by adapting the calibration quantile online. DtACI maintains coverage under arbitrary distribution shifts by adaptively reweighting a set of candidate adaptation rates based on their historical performance. It obtains a long-term coverage guarantee of the form in Eq. 1 in settings where the size of the distribution shift decays over time.
The DtACI algorithm seeks to compensate for distribution shift by changing the conformal quantiles $\alpha_{t}$ used to select the score threshold $q_{1-\alpha,t}$. The algorithm aims to minimize the difference between the optimal quantile $\alpha_{t}^{*}$ that would yield exact coverage at time $t$ and the estimated quantile $\alpha_{t}$. This difference is bounded in Theorem C.1. Observations in the DtACI algorithm are quantiles $\beta_{t}$, which are the lowest empirical quantiles of the dataset $\mathcal{D}_{t}$ that would have covered the true outcome at time $t$. Formally,
$\beta_{t}\coloneqq\sup\left\{\beta:Y_{t}\in\{y:s(X_{t},y;\theta_{t})\leq\hat{q}_{\beta,\mathcal{D}_{t}}\}\right\}$ (3)
DtACI maintains a set of $k$ candidate adaptation rates $\{\gamma_{i}\}_{i=1}^{k}$ and corresponding weights $\{w_{k,t}\}_{k=1}^{K}$. At each time step, the algorithm performs an exponential weights update on the weights based on the error incurred by each candidate rate. An averaged quantile $\bar{\alpha}_{t}$ is then computed using the weights, and the quantile threshold $q_{1-\bar{\alpha},\mathcal{D}_{t}}$ is updated accordingly. The algorithm is presented in Algorithm 3 (in Appendix B).
@@ 3 Related Work
Conformal prediction regions are used for applications in safety-critical systems, such as autonomous driving (Lindemann et al., 2023; Dixit et al., 2023), where it is crucial to maintain valid uncertainty estimates despite changing conditions. Efficient and adaptive conformal prediction methods can enhance the reliability of these systems by providing tighter uncertainty bounds that reflect the current data distribution. Dixit et al. (2023) demonstrate the utility of adaptive conformal prediction in autonomous driving scenarios, providing a framework for utilizing adaptive conformal methods in control. The ability to generate efficient prediction regions for control has been the subject of previous work, as the reduction in the volume of prediction regions directly impacts the conservativeness of control strategies (Lindemann et al., 2023; Tumu et al., 2024). However, these methods design static nonconformity scores without adaptation to distribution shifts, potentially leading to suboptimal performance in dynamic environments.
Conformal Prediction under Distribution Shift.
Two primary approaches have been proposed to extend conformal prediction to settings with distribution shift. The first approach is to modify the calibration set to better reflect the test distribution. This can be done by by reweighting calibration points based on their similarity to the test point (Barber et al., 2023; Tibshirani et al., 2019). These methods have been applied to multi-agent settings by Kuipers et al. (2024) and Binny & Dixit (2025).
The second approach is to adapt the quantile used for calibration over time, as in Adaptive Conformal Inference (ACI) (Gibbs & Candes, 2021) and other approaches like Multivalid Conformal Prediction (Bastani et al., 2022) and AgACI (Zaffran et al., 2022). These methods adjust the quantile based on recent coverage performance, allowing them to maintain coverage under certain types of distribution shift. DtACI allows for subsequent scores to be generated using different non-conformity scores, making it a natural fit for our framework.
Efficient Non-Conformity Scores.
The design of volume-efficient non-conformity scores has been the focus of recent work in conformal prediction for various environments (Kiyani et al., 2024; Gao et al., 2025), including control and robotics settings multi-dimensional outputs (Tumu et al., 2024, 2023; Cleaveland et al., 2024; Braun et al., 2025). Our approach extends this line of work by adapting the non-conformity score online to maintain efficiency under distribution shift.
@@ 4 AdaptNC Method
Our method, Adaptation of Non-Conformity scores (AdaptNC) is designed to maintain efficient and valid conformal prediction regions in settings with distribution shift. AdaptNC has three steps (1) Conformal Threshold Update (2) Score Parameter Adaptation, and (3) Replay. In the threshold update step, we choose the conformal quantile $q_{1-\bar{\alpha}_{t},\mathcal{D}_{t}}$ corresponding to the the current score function parameters. In the score parameter adaptation step, we use the current estimates of the rate of change of the data distribution to fit efficient non-conformity score parameters $\theta_{t}$ to the data. Finally, in the replay step, we account for the change in the distribution of the score functions by replaying recent observations, in order mitigating abrupt shifts in coverage. Due to computational intensity, steps (2) and (3) are only run every $t_{s}$ timesteps. The AdaptNC algorithm is presented pictorially in Fig. 1, and in detail in Algorithm 2 (in Appendix B).
@@
The aim of this step is to approximate the true conformal quantile $\alpha^{*}_{t}$, which is the quantile of the data such that $\Pr[Y_{t+1}\in\mathcal{C}_{t}(X_{t+1};\theta_{t},q_{\alpha^{*}_{t},\mathcal{D}_{t}})]\geq 1-\alpha$. The conformal threshold update step uses the framework of the DtACI algorithm (Algorithm 3) to update its estimate of the true conformal quantile $\alpha_{t}$ based on the observed scores. To enable coverage under unknown distribution shift, $k$ ACI (Gibbs & Candes, 2021) experts are maintained and dynamically reweighted. The procedure involves calculating an inverse quantile to transform the raw non-conformity score, expert reweighting, and finally expert update.
First, we save tuples $(X_{t},Y_{t},\hat{Y}_{t})$ in a history buffer $\mathcal{H}$. For each observation, $\beta_{t}$ is calculated according to Eq. 3. The dataset $\mathcal{D}_{t}$ used in the quantile calculation is a rolling window of the last $W$ timesteps. Based on the error between the predicted quantile for each expert $\alpha_{t}^{i}$ and the observed $\beta_{t}$, the weight $w_{t}^{i}$ is exponentially reweighted. When these weights are applied to the experts, we get the algorithm’s estimate of the current conformal quantile $\bar{\alpha}_{t}$. Finally, each expert is updated according to the expert learning rate $\gamma_{i}$ and the error indicator $\operatorname{\textbf{err}}_{t}^{i}$, which is $1$ if the observation is not covered by the $i$-th expert, and 0 otherwise. The complete DtACI algorithm is presented in Algorithm 3 (in Appendix B).
@@
The objective of this step is to find the parameters $\theta_{t}$ to best achieve the efficiency goal in Eq. 2. This problem is challenging because we do not know what the distribution will be in future timesteps. Fortunately, the expert weights $w_{t}^{i}$ and their expert learning rates $\gamma_{i}$ provide an estimate of how quickly the distribution is changing. In this step, we reweight the data in our history buffer and formulate an optimization problem to minimize the volume with respect to the current estimate of the data distribution.
First, we perform a reweighting of the data in the history buffer $\mathcal{H}$ based on our estimate of distribution shift. This estimate takes the form of the expert learning rates and weights $\gamma_{t},w_{t}^{i}$. Adaptive reweighting is important because it allows for the algorithm to “forget” past data based on the rates of change of the distribution when the distribution changes quickly, and utilize it for better estimates when the distribution is stable. For each historical timepoint, the weight $\omega_{t}$ is assigned based on Eq. 4, inducing the distribution $\mathcal{H}_{t}^{\omega}$.
$\bar{\omega_{t}}=\sum_{i=1}^{k}w_{t}^{i}(1-\gamma_{i})^{T-t+1},\quad\omega_{t}=\bar{\omega_{t}}/\sum_{j=1}^{|\mathcal{H}|}\bar{\omega_{j}}.$ (4)
Using this reweighted distribution, we can formulate the optimization problem below
@@
$\min_{\theta_{t}}\quad$ $\text{Volume}\left(\{y:s(X_{t},y;\theta_{t})\leq q_{1-\alpha,t}\}\right)$ (5a)
s.t. $\mathbb{P}_{(X_{t},Y_{t})\sim\mathcal{H}_{t}^{w}}(Y_{t}\in\hat{\mathcal{C}}_{t}(X_{t};\theta_{t},\mathcal{D}_{t}))\geq 1-\alpha$ (5b)
This optimization problem seeks to minimize the volume of the set covering $1-\alpha$ of the weight of the reweighted data distribution $\mathcal{H}_{t}^{w}$. This optimization can be performed using gradient-based methods or other suitable optimization techniques. In our approach, we use a variant of the method proposed in Tumu et al. (2024) to optimize the score parameters, while yielding convex shapes which are easy to use in downstream applications (Lindqvist et al., 2020). In this variant, we introduce a Monte-Carlo density estimation technique that avoids the computational complexity challenges in Tumu et al. (2024), and eliminate the clustering algorithm for simplicity.
This procedure takes place in two steps: (1) density estimation and (2) shape template optimization. Density estimation aims to identify the highest density regions of the reweighted data points in $\mathcal{H}$. Shape template optimization then fits convex shapes to cover these high-density regions, which are used to define the non-conformity score.
To identify the high-density region, we use the weighted sample points from our history buffer $\mathcal{H}$ to construct a Kernel Density Estimate (KDE) (Parzen, 1962). We identify the highest density regions that cover a proportion $1-\alpha$ of the total density mass through sampling a large number of points from the KDE and idenfifying their likelihoods. We select the top $1-\alpha$ proportion of these sampled points based on their likelihoods to form our set of high-density points $\mathcal{D}$. This yields a set of points $\hat{R}_{t}$ which comprise the high-density region. The complete algorithm is presented in Section B.4, along with a proof of asymptotic convergence.

### v2

# AdaptNC: Adaptive Nonconformity Scores for Conformal Prediction under Distribution Shift

@@ Abstract
Rigorous uncertainty quantification is essential for the safe deployment of autonomous systems in dynamic environments. Conformal Prediction (CP) provides a distribution-free framework for this task, yet its standard formulations rely on exchangeability assumptions that are violated by the distribution shifts inherent in real-world robotics. Existing online CP methods maintain target coverage by adaptively scaling the conformal threshold, but typically employ a static nonconformity score function. We show that this fixed geometry leads to highly conservative, volume-inefficient prediction regions when environments undergo structural shifts. To address this, we propose AdaptNC, a framework for the joint online adaptation of both the nonconformity score parameters and the conformal threshold. AdaptNC leverages an adaptive reweighting scheme to optimize score functions, and introduces a replay buffer mechanism to mitigate the coverage instability that occurs during score transitions. We evaluate AdaptNC on diverse robotic benchmarks involving multi-agent policy changes, environmental changes and sensor degradation. Our results demonstrate that AdaptNC significantly reduces prediction region volume compared to state-of-the-art threshold-only baselines while maintaining target coverage levels.
@@ 1 Introduction
Autonomous systems operating in real-world environments are subject to constant change. These changes induce distribution shift, driven by changing environmental conditions, such wind or weather, system degradation due to damage or wear and tear, or through multi-agent interactions, which can all degrade the performance of predictive models. For safety-critical applications such as autonomous driving and robotic navigation, identifying these shifts and providing robust, uncertainty-aware predictions is not just a performance requirement; it is a safety necessity [7]. Distribution shift in robotics does not only change the magnitude or frequency of errors, it can change the structure of the errors. For example, tire damage in a ground vehicle or strong winds affecting a drone can induce directionally biased prediction errors. Because prediction regions are often used to define where robots should slow down, avoid, or plan conservatively [18], these structural changes directly affect the usefulness of uncertainty estimates. Thus, prediction regions should not only adapt in size, but also in shape, as environmental and system conditions evolve.
Uncertainty quantification has been a focus of much recent work, and conformal prediction (CP) [28, 17], has emerged as a popular method for its flexibility to be applied to predictors of all types, its statistical grounding, and its ease of use for the practitioner. CP constructs prediction sets which contain the true label with high ($1-\alpha$) probability, under the assumption of exchangeability between the calibration and test distributions. In CP, the threshold controls the size of the prediction region, while the non-conformity score function controls its geometry. Online adaptation has largely focused on the former, leaving the latter fixed even when the structure of the data changes.
Recent online CP methods adapt the conformal threshold to maintain coverage under distribution shift[12, 11, 29]. However, because they keep the non-conformity score fixed, they can only resize prediction regions rather than reshape them. This limitation is important in robotic settings, where shifts such as sensor degradation or policy changes can alter the geometry of prediction errors. Separately, previous work has addressed the design of non-conformity score functions[15, 10], including control and robotics settings with multi-dimensional outputs [25, 26, 9, 6], however, this work does not address the adaptation of the non-conformity score in response to distribution shift.
@@
Unlike existing methods that only tune the threshold, AdaptNC optimizes the underlying score function to ensure that prediction regions remain tight and informative even as the data distribution evolves. AdaptNC also shifts the conformal threshold. In order to provide principled score function adaptation, we introduce an adaptive reweighting scheme that leverages DtACI expert weights to prioritize the most relevant historical observations during score optimization. The joint adaptation of the score function and threshold introduces the phenomenon of “coverage shock": the errors in coverage caused by evaluating old threshold values under a new score function. To mitigate “coverage shock” we employ a counterfactual replay mechanism to recalibrate the conformal threshold following a score update. Across three robotic settings with structural distribution shifts, AdaptNC maintains target coverage while producing smaller prediction regions than threshold-adaptation baselines, demonstrating that score adaptation improves efficiency without sacrificing empirical coverage.
@@
Contributions: This paper addresses the challenge of constructing efficient, uncertainty-aware prediction regions under distribution shift by contributing:
@@
    Online Score Function and Threshold Adaptation: We present a method for the online optimization of non-conformity scores and their corresponding conformal thresholds that minimizes prediction region volume while enabling adaptation to shifting distributions.
@@
    Dynamic Weighting and Replay Buffer: We develop a weighting mechanism for data based on the rate of distribution shift, and a replay mechanism to maintain coverage stability during score parameter transitions.
@@
    Empirical Validation on Robotic Benchmarks: We evaluate AdaptNC on robotic tasks involving structural distribution shifts, including multi-agent policy changes and sensor degradation. We demonstrate that AdaptNC reduces prediction volume compared to state-of-the-art online CP baselines while maintaining target coverage levels.
Furthermore, we emphasize that AdaptNC is not a direct composition of prior adaptive conformal prediction and score-adaptation methods. Instead, it formulates a new setting involving coupled adaptation, characterizes the instabilities arising from a naive integration, and introduces the mechanisms necessary to ensure stable and effective joint adaptation. Additional discussion is provided in Appendix B.1.
@@ 2 Problem Setup and Preliminaries
2.1 Online Prediction Setting
We consider an online $\tau$-step prediction setting. At each time step $t=1,\ldots,T$, the predictor observes features $X_{t}\in\mathcal{X}$ and produces a $\tau$-step prediction $\hat{Y}_{t}=h(X_{t})=\{\hat{y}_{t+1},\ldots,\hat{y}_{t+\tau}\}\in\mathcal{Y}^{\tau}$. The future outcome $Y_{t}\in\mathcal{Y}^{\tau}$ is then observed. The joint distribution of $(X_{t},Y_{t})$ may change over time.
We seek to build a prediction region $\mathcal{C}_{t}$ that contains the future outcome $Y_{t}$ with probability $1-\alpha$, where $\alpha\in(0,1)$ is the target miscoverage rate. The prediction regions are constructed using a non-conformity score function $s(X_{t},y;\theta_{t})$ where $\theta_{t}$ are parameters of the score function at time $t$, as well as a conformal threshold, $q_{t}$. We define $q_{\mathcal{D}_{t}}(\delta)$ to be the empirical $\delta\in[0,1]$ quantile of the score dataset $\mathcal{D}_{t}$, and $q_{t}=q_{\mathcal{D}_{t}}(\delta_{t})$. The prediction region is the set of points with a non-conformity score less than the conformal quantile: $\mathcal{C}_{t}(X_{t};\theta_{t},q_{t})\coloneqq\{y\in\mathcal{Y}^{\tau}:s(X_{t},y;\theta_{t})\leq q_{t}\}$. The threshold $q_{t}$ controls the size of the prediction region, while the score parameters $\theta_{t}$ control its geometry.
Desiderata:
We pursue the objectives of coverage that approaches the target, and efficiency of the prediction region under settings with distribution shift.
Coverage: Under distribution shift, exact finite-sample coverage at every time step is generally not attainable. Instead, we seek the asymptotic miscoverage rate $\alpha$, in settings where the magnitude of distribution shift decays over time. We define the error indicator $\operatorname{\textbf{err}}_{t}=\mathds{1}\{Y_{t}\notin\mathcal{C}_{t}(X_{t};\theta_{t},q_{t})\}$, and we target:
@@ 2.1 Online Prediction Setting
Efficiency: We would also like to ensure that the prediction regions $\mathcal{C}_{t}$ are small and informative to downstream algorithms. Subject to our coverage desideratum, our objective is to minimize the volume of the prediction set $\mathcal{C}_{t}$. We formulate the optimization problem:
$\min_{\theta_{t}}\quad$ $\text{Vol}(\mathcal{C}_{t}(X_{t+1};\theta_{t},q_{t})),\quad\text{s.t.}~~\Pr[Y_{t+1}\in\mathcal{C}_{t}(X_{t+1};\theta_{t},q_{t})]\geq 1-\alpha.$ (2a)
In the online setting, the future distribution is unknown, so this problem serves as an idealized objective: AdaptNC seeks to reduce prediction-region volume while maintaining long-run coverage.
@@ 2.2 Preliminaries
Conformal Prediction: Conformal prediction (CP) and its split variants provide distribution-free, finite-sample prediction regions which guarantee ground truth coverage with $1-\alpha$ probability under the assumption of exchangeable calibration and test data [28, 17, 3]. Under distribution shift, this exchangeability assumption may fail, motivating online methods that adapt the threshold over time.
Online Threshold Adaptation: Online conformal prediction methods adapt the threshold to maintain long-run coverage in settings where distribution shift occurs. DtACI [11] targets long-run coverage under arbitrary distribution shifts by adaptively reweighting a set of candidate adaptation rates based on their historical performance. It obtains a long-term coverage guarantee of the form in Eq.˜1 in settings where the size of the distribution shift decays over time.
The DtACI algorithm seeks to compensate for distribution shift by changing the adaptive miscoverage level $\alpha_{t}$ used to select the score threshold $q_{t}=q_{\mathcal{D}_{t}}(1-\alpha_{t})$. The algorithm seeks to find $\alpha_{t}$ that minimizes the regret to the hindsight optimal quantile $\alpha^{*}_{t}$. This difference is bounded in Theorem˜D.1. Observations in the algorithm are quantiles $\beta_{t}$, which are the lowest empirical quantiles of the dataset $\mathcal{D}_{t}$ that would have covered the true outcome at time $t$. Formally,
$\beta_{t}\coloneqq\inf\left\{\beta:Y_{t}\in\{y:s(X_{t},y;\theta_{t})\leq\hat{q}_{\mathcal{D}_{t}}(\beta)\}\right\}$ (3)
To achieve this, DtACI maintains a set of $k$ candidate adaptation rates $\{\gamma_{i}\}_{i=1}^{k}$ and corresponding weights $\{w_{t}^{i}\}_{i=1}^{k}$. At each time step, the algorithm performs an exponential weights update on the weights based on the error incurred by each candidate rate. An averaged quantile $\bar{\alpha}_{t}$ is then computed using the weights, and the quantile threshold $q_{\mathcal{D}_{t}}(1-\bar{\alpha}_{t})$ is updated accordingly. The algorithm is presented in Algorithm˜3 (in Appendix C).
@@ 3 Related Work
Online Threshold Adaptation.
Online conformal prediction methods address distribution shift primarily by adapting the conformal threshold or calibration quantile over time. Adaptive Conformal Inference (ACI) updates the quantile based on recent coverage errors [12], while related methods such as AgACI and multivalid conformal prediction provide alternative mechanisms for maintaining coverage under non-stationarity [29, 5]. DtACI further advances this line of work by adaptively weighting multiple candidate adaptation rates and providing long-term coverage regret guarantees [11]. However, these online CP methods adapt the threshold while leaving the non-conformity score fixed. As a result, they can resize prediction regions under distribution shift, but they cannot reshape them when the geometry of prediction errors changes. Other approaches address distribution shift by reweighting calibration points to better match the test distribution under covariate shift [24, 4]. These methods modify the effective calibration distribution, and have been used in reinforcement learning settings [23, 16] but they do not adapt the non-conformity score online.
Efficient and Shape-Adaptive Non-Conformity Scores.
A separate line of work studies how the choice of non-conformity score affects the shape and volume of conformal prediction regions. Recent methods design volume-efficient scores for general prediction settings [15, 10], as well as for control and robotics problems with multi-dimensional outputs [25, 26, 9, 6]. These methods show that intentional design of the non-conformity score function can substantially improve prediction-region efficiency and downstream usefulness. However, they typically design or optimize the score for a fixed data-generating process. They do not consider online score adaptation under distribution shift, where the geometry of prediction errors may change over time.
@@ 4 AdaptNC Method
Our method, Adaptation of Non-Conformity Scores (AdaptNC), extends online conformal prediction by jointly adapting the conformal threshold and non-conformity score function (NCSF). (1) Threshold adaptation maintains long-run coverage, (2) Score Parameter Adaptation improves efficiency under distribution shift, and (3) Replay recalibrates the threshold after score updates to mitigate coverage shock. At each timestep, the conformal threshold is updated using the quantile $q_{\mathcal{D}_{t}}(1-\bar{\alpha}_{t})$ based on the current NCSF and newly observed data. Score parameters $\theta_{t}$ are periodically updated using estimates of distribution shift to obtain more efficient regions, and replay rescores recent observations under the updated NCSF ccount for the induced change in the score distribution . Due to computational intensity, steps (2) and (3) are performed every $t_{s}$ timesteps, while (1) is updated at every timestep. An overview is shown in Fig.˜1, with full details in Algorithm˜2 (Appendix C).
The data observed is saved as tuples $(X_{t},Y_{t},\hat{Y}_{t})$ in a history buffer $\mathcal{H}$. For each observation, $\beta_{t}$ is calculated according to Eq.˜3. The dataset $\mathcal{D}_{t}$ used in the quantile calculation is a rolling window of the last $W$ timesteps.
@@
At every timestep, AdaptNC uses DtACI (Algorithm˜3) to adapt the conformal threshold online. The goal of this step is to select an adaptive miscoverage level $\bar{\alpha}_{t}$ and the corresponding quantile $q_{t}=q_{\mathcal{D}_{t}}(1-\bar{\alpha}_{t})$, where $\mathcal{D}_{t}$ is a rolling window of the last $W$ timesteps. To enable coverage under unknown distribution shift, $k$ ACI [12] experts are maintained and dynamically reweighted. The procedure involves calculating an inverse quantile to transform the raw non-conformity score, expert reweighting, and finally expert update. This procedure aims to estimate $\alpha^{*}_{t}$, the optimal value for the miscoverage rate that would yield the correct coverage.
Based on the error between the predicted quantile for each expert $\alpha_{t}^{i}$ and the observed $\beta_{t}$, the weight $w_{t}^{i}$ is exponentially reweighted. When these weights are applied to the experts, we get the algorithm’s estimate of the current conformal quantile $\bar{\alpha}_{t}$. Finally, each expert is updated according to the expert learning rate $\gamma_{i}$ and the error indicator $\operatorname{\textbf{err}}_{t}^{i}$, which is $1$ if the observation is not covered by the $i$-th expert, and 0 otherwise. The complete DtACI algorithm is presented in Algorithm˜3 (in Appendix C).
@@
Threshold adaptation can maintain coverage under distribution shift, but it can only resize prediction regions for a fixed score function. To adapt the geometry of the prediction region, AdaptNC periodically updates the score parameters $\theta_{t}$. The objective of this step is to find score parameters that reduce prediction-region volume while preserving the desired coverage behavior. Since the future distribution is unknown, AdaptNC approximates this objective using a relevance-weighted history buffer, where the relevance weights are derived from the current DtACI expert weights $w_{t}^{i}$ and adaptation rates $\gamma_{i}$.
Let $\mathcal{H}_{t}$ denote the history buffer at time $t$. For each historical observation indexed by $j\leq t$, AdaptNC assigns a relevance weight using the DtACI expert mixture:
$\bar{\omega}_{j,t}=\sum_{i=1}^{k}w_{t}^{i}(1-\gamma_{i})^{t-j},\qquad\omega_{j,t}=\frac{\bar{\omega}_{j,t}}{\sum_{\ell\in\mathcal{H}_{t}}\bar{\omega}_{\ell,t}}$ (4)
This weighting scheme induces a weighted empirical distribution $\mathcal{H}_{t}^{\omega}$. Experts with larger adaptation rates $\gamma_{i}$ emphasize recent observations, while experts with smaller adaptation rates retain longer memory. Therefore, the DtACI mixture provides an adaptive estimate of how much historical data should influence score adaptation.
Using the weighted history distribution $\mathcal{H}_{t}^{\omega}$, AdaptNC seeks score parameters that minimize prediction-region volume subject to weighted empirical coverage. This gives the objective
@@
$\min_{\theta}\quad\mathbb{E}_{X\sim\mathcal{H}_{t}^{\omega}}\left[\mathrm{Vol}\left(\mathcal{C}(X;\theta,q_{\theta})\right)\right]\quad\text{s.t.}~~\mathbb{P}_{(X,Y)\sim\mathcal{H}_{t}^{\omega}}\left[Y\in\mathcal{C}(X;\theta,q_{\theta})\right]\geq 1-\alpha,$ (5a)
where $q_{\theta}$ denotes the threshold induced by the score function with parameters $\theta$ on the weighted history. In practice, AdaptNC approximates this objective by estimating a high-density region of the weighted residual distribution and fitting a convex score template to that region.
To estimate the high-density region, AdaptNC forms residual samples $Y_{j}-h(X_{j})$ from the weighted history buffer and constructs a weighted kernel density estimate (KDE) [20]. We then sample candidate residuals from the KDE, evaluate their estimated density, and retain the highest-density samples whose empirical mass is approximately $1-\alpha$. We denote this retained set by $\hat{R}_{t}$. This Monte Carlo procedure provides an empirical approximation of the high-density residual region under the current weighted data distribution; the full algorithmic details are given in Appendix C.4.